find the center,foci,vertices, endpoints of conjugate axis.determine the equation of the asymptotes and sketsch the graph



1. (y+6)²/25-(x-4)²/39=1



2. 9x²+126x-16y²-96y+153=0


1
Expert's answer
2021-11-25T23:48:20-0500

1.


"\dfrac{(y+6)^2}{25}-\dfrac{(x-4)^2}{39}=1"

Center: "(4, -6)"

Vertices: "(4, -11), (4, -1)"

"c^2=a^2+b^2=25+39=64=>c=8"

Foci: "(4,-14), (4, 2)"

Endpoints of conjugate axis: "(4-\sqrt{39}, -6), (4+\sqrt{39}, -6)"


First asymptote: "y=-\dfrac{5\sqrt{39}}{39}(x-4)-6"


Second asymptote: "y=\dfrac{5\sqrt{39}}{39}(x-4)-6"




2.


"9x^2+126x-16y^2-96y+153=0"

"9(x^2+14x+49)-16(y^2+6y+9)=144"

"\dfrac{(x+7)^2}{16}-\dfrac{(y+3)^2}{9}=1"

Center: "(-7, -3)"

Vertices: "(-11, -3), (-3, -3)"

"c^2=a^2+b^2=16+9=25=>c=5"

Foci: "(-12,-3), (-2,-3)"

Endpoints of conjugate axis: "(-7, -6), (-7,0)"


First asymptote: "y=-\dfrac{3}{4}x-\dfrac{33}{4}"


Second asymptote: "y=\dfrac{3}{4}x+\dfrac{9}{4}"